Stan can run 1 miles per hour, wheras Jack can run only 8 miles per hour. If they start at the same time from the same point and run in opposite directions, how far apart (to the nearest mile) will they be after 10 minutes?
A) 1 mile
B) 2 miles
C) 3 miles
D) 4 miles
E) 5 miles

Answer by Alan3354(69443)

Stan can run 1 miles per hour, wheras Jack can run only 8 miles per hour. If they start at the same time from the same point and run in opposite directions, how far apart (to the nearest mile) will they be after 10 minutes?
-----------
Speeds are added when going in opposite directions.
1 + 8 = 9 mi/hr
10 minutes = 1/6 hour
---
d = r*t = 9*(1/6)
d = 1.5 miles